Solve the following system:
x-3y+4z=7
2x+3y-6z=5
3x-2y-z=14

Answer by stanbon(48568)

Solve the following system:
x-3y+4z=7
2x+3y-6z=5
3x-2y-z=14
--------------
x = 4
y = -1
z = 0
==========
Cheers,
Stan H.